# Segment Anything Model (SAM) for Radiation Oncology

### v1

@@ Abstract
In this study, we evaluate the performance of the Segment Anything Model (SAM) model in clinical radiotherapy. We collected real clinical cases from four regions at the Mayo Clinic: prostate, lung, gastrointestinal, and head & neck, which are typical treatment sites in radiation oncology. For each case, we selected the OARs of concern in radiotherapy planning and compared the Dice and Jaccard outcomes between clinical manual delineation, automatic segmentation using SAM’s "segment anything" mode, and automatic segmentation using SAM with box prompt. Our results indicate that SAM performs better in automatic segmentation for the prostate and lung regions, while its performance in the gastrointestinal and head & neck regions was relatively inferior. When considering the size of the organ and the clarity of its boundary, SAM displays better performance for larger organs with clear boundaries, such as the lung and liver, and worse for smaller organs with unclear boundaries, like the parotid and cochlea. These findings align with the generally accepted variations in difficulty level associated with manual delineation of different organs at different sites in clinical radiotherapy. Given that SAM, a single trained model, could handle the delineation of OARs in four regions, these results also demonstrate SAM’s robust generalization capabilities in automatic segmentation for radiotherapy, i.e., achieving delineation of different radiotherapy OARs using a generic automatic segmentation model. SAM’s generalization capabilities across different regions make it technically feasible to develop a generic model for automatic segmentation in radiotherapy.
@@ 1 Introduction
Recent advancements in natural language processing have led to large language models (LLMs) that can generalize to new domains with little data [1, 2]. Models like GPT-3 [3], ChatGPT [2], GPT-4 [4] and Google’s PaLM-2 [5] have revolutionized NLP, enabling AI systems that can perform a wide range of language tasks [6, 7, 8, 9, 10, 11, 12, 13]. Their success has inspired interest in building similar "foundation models" for computer vision [14, 15].
@@
In response, Meta’s Segment Anything Model (SAM) [16] was proposed as a generalized and promptable model for image segmentation. SAM is trained on over 1 billion masks and can generate segmentation masks for any object based on prompts. SAM shows strong performance on natural images and may enable zero-shot learning for new objects without retraining. SAM’s capability suggests it could enhance interactive medical image segmentation where physicians provide guidance to generate accurate delineations. While SAM’s performance on natural images is impressive, medical images possess unique challenges. SAM was not designed specifically for medical images and may struggle with their complexity [15].
@@
Since SAM was proposed, various foundation models have emerged. In image editing, the inpaint anything (IA) framework [17] has successfully integrated SAM with state-of-the-art image inpainters [18] and AI-generated content (AIGC) models [19]. This integration has resulted in a powerful pipeline capable of addressing various challenges in inpainting-related tasks. Another notable contribution is the "edit everything" approach [20], which follows a similar pipeline to IA. In this case, SAM is utilized to segment the input image without any prompts, and subsequently, a source prompt is employed to guide CLIP for image editing using simple text instructions. Furthermore, SAM has also found applications in style transfer. The Any-to-Any Style Transfer [21] leverages SAM’s promotable region selection capability to facilitate effective style transfer between different images. Additionally, SAM’s versatility extends to object detection tasks. Giannakis et al. [22] propose a universal crater detection scheme utilizing the zero-shot generalization of SAM, enabling the detection of unfamiliar objects. In addition to its various applications in natural images, SAM has been evaluated across diverse real-world segmentation scenarios [23] involving different types of images, such as medical images in the healthcare domain.
@@
In recent years, an escalated incidence of cancer has been observed, with approximately half of these cases necessitating radiation therapy (RT), as indicated by evidence-based assessments [24]. Although efficacious in eradicating tumor cells, RT concurrently poses the risk of damaging normal tissues, potentially leading to an array of complications. It is therefore of paramount importance to accurately delineate OARs proximal to the tumor on simulated computed tomography (CT) images prior to formulating the treatment plan. To minimize tissue damage, the dose to these organs should be reduced during the plan optimization phase, guided by the accurately delineated OARs[25, 26, 27, 28, 29]. Currently, the process of OAR delineation is manual and is executed by a radiation oncologist or a dosimetrist. It involves an in-depth analysis of CT images on a slice-by-slice basis, a procedure that is both labor-intensive and time-consuming[30, 31]. The burgeoning demand for RT inevitably extends patient waiting times, a factor shown to adversely impact local control and prognosis due to tumor proliferation[32]. For adaptive RT, which necessitates frequent delineation of OARs and target volumes on updated CT images along with alterations to treatment plans during the course of treatment, the need for expeditious and accurate delineation is even more pronounced. Furthermore, imprecise delineation may lead to dose miscalculations and unintended complications. Successful radiation therapy necessitates high geometric and dosimetric precision[33, 34, 35, 36, 37]. The rising trend of pencil beam scanning proton therapy has afforded us the ability to more accurately shape dose distributions to align with the treatment target, whilst minimizing the exposure to organs at risk (OAR)[38, 39, 40, 41, 42]. Nonetheless, the sharp dose gradients facilitated by these methods could potentially pose a risk if the delineation of structures is imprecise.Therefore, the rapid and precise delineation of OARs becomes vital, especially in regions where there is a dearth of experienced radiation oncologists.
@@
An automatic segmentation (AS) algorithm can potentially address these challenges, thereby considerably enhancing RT efficiency [43]. This algorithm is capable of accurately delineating OARs within a reasonable timeframe, thereby becoming a crucial tool in the RT process. Numerous auto-segmentation methodologies have been proposed, including deformable image registration (DIR), atlas-based auto-segmentation, and the more recent deep learning-based segmentation (DLS) [44]. Although both DIR and atlas-based auto-segmentation have seen extensive implementation, their clinical utility is compromised due to limitations in accuracy and efficiency [45, 46]. The focus of auto-segmentation research has recently shifted towards artificial intelligence (AI) methods, with a particular emphasis on those underpinned by deep learning (DL) [43, 47]. Over the past few years, the number of studies and clinical applications exploring DL-based segmentation of OARs in RT has proliferated, significantly improving the efficiency of auto-delineation in RT processes [48, 49, 50]. However, clinical feedback suggests a prominent issue with the current models: a lack of generalizability [51, 52]. These models often necessitate the training of a unique model for each RT site. Given the variety of RT sites, a multitude of independent auto-segmentation models need to be trained, each requiring a considerable investment of time and a large volume of site-specific training data. In practical clinical scenarios, the determination of the site to be delineated often requires the invocation of a site-specific model. The complexity inherent in such scenarios, including mixed-site images and images from rare sites, can significantly degrade the precision of existing auto-segmentation models, and in some cases, cause errors. Thus, the development of a universal auto-delineation model, capable of handling multiple sites simultaneously, could significantly augment the precision and efficiency of auto-segmentation in RT processes, particularly in delineating OARs. An ideal solution would rely on limited data, generalize across modalities and anatomies, and minimize human effort. SAM demonstrates such potential but its performance on medical images remains unclear, especially for radiotherapy use cases [53, 15].
@@
We evaluate SAM’s ability to perform zero-shot segmentation of medical images from multiple anatomical sites from clinical radiation oncology with CT images (prostate, lung, gastrointestinal, head&neck). We assess SAM in "segment everything" mode where it generates masks for all objects and "manual box prompt" mode where users indicate regions of interest. Dice coefficient and Jaccard index measure the spatial overlap between SAM’s predictions and ground truth clinical delineations.
@@
Our results provide a broad analysis of SAM for radiation oncology. SAM shows promise but performance varies significantly based on the region and modality. SAM excels at segmenting large, well-defined organs given unambiguous prompts but struggles at complex anatomies with ambiguous prompts, especially smaller organs. Performance metrics suggest general trends but fuller qualitative analysis is needed to determine practical utility, and there is much room for adapting and improving SAM for radiation oncology applications.
@@ 2.1 Segment Anything Model
The Segment Anything project proposed by Meta is a groundbreaking initiative that aims to democratize the world of image segmentation, a vital task in computer vision [16]. It consists of two key components: a substantial dataset for image segmentation, and the SAM, a promotable foundation model. Taking inspiration from the world of NLP, the project creates a vast dataset and a segmentation model, both open-sourced, thus opening up enormous possibilities. SAM, which can run in real-time in browsers, offers a highly automated image segmentation approach that requires minimal human intervention. It’s a deep learning model, trained on over 1 billion masks in 11 million images, which can cut out almost anything from an image. Unlike traditional models that require specialized training, SAM is generalizable and can respond to user prompts about specific areas to segment.
This model includes three essential components: an image encoder, a prompt encoder, and a mask decoder. Input images pass through an image encoder to produce an embedding, and the model can accept prompts as points, boxes, or rough masks. For more nuanced prompting, the authors are working on a version of SAM that accepts text input, similar to language models. The resultant segmented image comes with multiple valid masks and a confidence score, signifying the segmentation’s accuracy. SAM’s capabilities could revolutionize various fields, from augmenting reality through precise object identification to biomedical image segmentation, aiding cell microscopy analysis [54]. It can also be integrated with diffusion-based image generation models for efficient image inpainting and be utilized for generating semantic segmentation datasets, making it an exciting development in the realm of artificial intelligence and computer vision [55].
@@ 2.2 SAM for Medical Imaging
Since SAM was proposed as an innovative framework for image segmentation, its application within the realm of medical imaging has increasingly become a topic of interest within the field [15] [56]. SAM has been tested under the ’everything’ mode for segmenting lesion regions across a range of anatomical structures, such as the brain, lung, and liver, and imaging modalities like Computerized Tomography (CT) and Magnetic Resonance Imaging (MRI). The experimental findings suggest that while SAM exhibits relative proficiency in segmenting organs with distinct boundaries, it encounters difficulties when trying to accurately identify lesions with irregular shapes and contours [57]. Another study drew comparisons between the Segmentation Anything Model (SAM) and the Brain Extraction Tool (BET) from the FMRIB Software Library in brain extraction tasks, and the quantitative analysis revealed superior segmentation performance by SAM compared to BET, thus illustrating SAM’s promising potential for use in brain segmentation tasks [58]. SAM was also applied to digital pathology segmentation tasks, which encompassed the segmentation of tumor and non-tumor tissue, as well as cell nuclei, on whole-slide images. The findings indicated that SAM exhibits excellent performance when segmenting large, interconnected objects. However, its reliability wavers when tasked with dense instance object segmentation, even when provided with all the target boxes [59]. SAM for polyp segmentation tasks using five benchmark datasets, all under the ’everything’ setting has been reported. The findings indicate that while SAM has the capacity to accurately segment polyps in certain instances, it significantly lags behind the performance of leading-edge methods, suggesting substantial room for improvement [60]. Extensive experimentation on multiple public datasets shows that the zero-shot segmentation capabilities of SAM are not sufficiently robust for direct application in medical image segmentation [15] [61] [54].
@@ 2.3 Cancer Segmentation
The field of cancer segmentation in medical imaging has seen considerable advancements in recent years, motivated by the potential to improve the accuracy of cancer detection, diagnosis, and treatment planning. The goal of cancer segmentation is to accurately delineate the boundaries of cancerous regions within medical images.
@@
The recent explosion of deep learning techniques has revolutionized the field of cancer segmentation [65]. For instance, U-Net [66], a type of CNN, has been widely used for biomedical image segmentation tasks including lung cancer segmentation from CT scans and brain tumor segmentation from MRI images [66]. Developing fully automated systems for cancer segmentation is a significant trend in recent research [67]. For example, a recent study explored the possibility of using neural architecture search (NAS) to find optimal parameters for U-Net based models and achieved significant improvement over baselines [68]. These systems aim to minimize the need for manual intervention, thus reducing the time taken for diagnosis and the potential for human error.
The use of SAM for cancer segmentation has seen a significant increase since it was proposed. With point or bounding box prompts, SAM achieves competitive results comparable to supervised training models like U-net in the areas of liver tumor, breast tumor, and colon polyp segmentation [69, 70, 71]. By fine-tuning the image encoder, prompt encoder, and mask decoder using the same labeled data, SAM can even outperform state-of-the-art results in skin cancer segmentation [72]. These studies demonstrate that the pre-trained SAM is efficient and effective for computer-aided rapid cancer examination and diagnosis. Furthermore, it can also serve as a versatile foundation model for developing domain-specific models fine-tuned for targeted cancer segmentation.
@@ 3.1 Datasets
This study has received approval from the Institutional Review Board (IRB) at Mayo Clinic. Fig. 1 illustrates the overall framework of our work. We collected case images from the four most common sites in clinical radiotherapy, namely the prostate, digestive system, lungs, and head and neck. We gathered 20 cases from each site, totaling 80 cases. From the perspective of clinical radiotherapy delineation, we divided the cases into two groups: the simple group includes the prostate and lungs, while the difficult group encompasses the digestive system and head and neck. As SAM currently supports 2D delineation, for a fair comparison, we evaluated the delineation of five typical 2D slices extracted from each image that best represented the anatomical information. Based on the organs that require attention during the formulation of the radiotherapy plan, we selected regions of interest for delineation comparison for different case sites, following the recommendations of the Radiation Therapy Oncology Group (RTOG).
For the prostate region, we selected the following regions of interest for evaluation: prostate, bladder, left femoral head, right femoral head, and rectum. For the lungs, we evaluated the following areas: left lung, right lung, heart, spinal cord, and esophagus. For the digestive system, we assessed the liver, stomach, left kidney, right kidney, spinal cord, large bowel, and small bowel. For the head and neck, we evaluated the brain, left parotid, right parotid, spinal cord, mandible, left cochlea, and right cochlea. All manual delineations were performed by highly experienced radiotherapy physicians and meet the RTOG delineation standards, and are used in the formulation of clinical radiotherapy plans.

### v2

@@ Abstract
In this study, we evaluate the performance of the Segment Anything Model (SAM) in clinical radiotherapy. We collected clinical cases from four disease sites at Mayo Clinic: prostate, lung, gastrointestinal, and head & neck, which are major treatment sites in radiation oncology. For each case, we selected the organs-at-risk (OARs) in radiotherapy planning. We then compared both the Dice coefficients and Jaccard indices derived from three distinct methods: clinical manual delineation (considered to be the ground truth), automatic segmentation using SAM’s ’segment anything’ mode, and automatic segmentation using SAM’s ’box prompt’ mode. Our results indicate that SAM’s segment anything mode can achieve clinically acceptable segmentation results in most OARs with Dice scores higher than 0.7. SAM’s box prompt mode further improves the Dice scores by 0.1$\sim$0.5. The results show that SAM performs better in automatic segmentation for the prostate and lung sites, while its performance lags behind in the gastrointestinal and head & neck sites. When considering the size of the organ and the clarity of its boundary, SAM displays better performance for large organs with clear boundaries, such as the lung and liver, and worse for smaller organs with unclear boundaries, like the parotid and cochlea. These findings align with the generally accepted variations in difficulty level associated with the manual delineation of different organs at different sites in clinical radiotherapy. Given that SAM, a model pre-trained purely based on natural images, could handle the delineation of OARs from medical images with clinically acceptable accuracy, these results highly demonstrate SAM’s robust generalization capabilities with consistent accuracy in automatic segmentation for radiotherapy, i.e., achieving delineation of different OARs of different sites using a generic automatic segmentation model. SAM’s generalization capabilities across different disease sites make it technically feasible to develop a generic model for automatic segmentation in radiotherapy. Further research to enhance SAM to support 3D segmentation and multiple modalities is warranted.
@@ 1 Introduction
Recent advancements in natural language processing (NLP) have led to large language models (LLMs) that can generalize to new domains with little training data [1, 2]. Models such as GPT-3 [3], ChatGPT [2], GPT-4 [4], and Google’s PaLM-2 [5] have revolutionized NLP. LLMs enable artificial intelligence (AI) systems that can perform a wide range of language tasks [6, 7, 8, 9, 10, 11, 12, 13] in diverse domains [14, 15, 7]. Their significant success has inspired interest in building similar "foundation models" for computer vision [16, 17, 1].
@@
In response, Meta’s Segment Anything Model (SAM) [18] was proposed as a generalized and promptable model for image segmentation. SAM is trained on over 1 billion masks in 11 million natural images and can generate segmentation masks for any object based on prompts. SAM shows strong performance on natural images and may enable zero-shot learning for new objects without retraining. SAM’s capability suggests that it could enhance interactive medical image segmentation where physicians provide guidance to generate accurate delineations. While SAM’s performance on natural images is impressive, medical images possess unique challenges. SAM was not designed specifically for medical images and may struggle with their complexity [17].
@@
Since SAM was proposed, various foundation models have emerged. In image editing, the inpaint anything (IA) framework [19] has successfully integrated SAM with state-of-the-art image inpainters [20] and AI-generated content (AIGC) models [21]. This integration has resulted in a powerful pipeline capable of addressing various challenges in inpainting-related tasks. Another notable contribution is the "edit everything" approach [22], which follows a similar pipeline to IA. In this case, SAM is utilized to segment the input image without any prompts, and subsequently, a source prompt is employed to guide CLIP (Contrastive Language–Image Pre-training) [23] for image editing using simple text instructions. Furthermore, SAM has also found applications in style transfer. The Any-to-Any Style Transfer [24] leverages SAM’s promotable region selection capability to facilitate effective style transfer between different images. Additionally, SAM’s versatility extends to object detection tasks. Giannakis et al. [25] propose a universal crater detection scheme utilizing the zero-shot generalization of SAM, enabling the detection of unfamiliar objects. In addition to its various applications in natural images, SAM has been evaluated across diverse real-world segmentation scenarios [26] involving different types of images, such as medical images in the healthcare domain.
@@
In recent years, an escalated incidence of cancer has been observed, with approximately half of these cases necessitating radiation therapy (RT), as indicated by evidence-based assessments [27]. Although efficacious in eradicating tumor cells, RT concurrently poses the risk of damaging nearby normal tissues, potentially leading to an array of complications. It is therefore of paramount importance to accurately delineate OARs proximal to the tumor on simulated computed tomography (CT) images prior to formulating the treatment plan. To minimize tissue damage, the dose to these OARs should be minimized as much as possible during the treatment planning, guided by the accurately delineated OARs [28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Currently, the process of OAR delineation is manual and is executed by a radiation oncologist or a dosimetrist. It involves an in-depth analysis of CT images on a slice-by-slice basis, a procedure that is both labor-intensive and time-consuming [38, 39]. The burgeoning OAR delineation demand for RT inevitably extends patients’ waiting times, a factor shown to adversely impact tumor local control and prognosis due to tumor proliferation [40]. For adaptive RT, which necessitates frequent delineation of OARs and target volumes on verification CT images along with alterations to treatment plans during the course of treatment, the need for expeditious and accurate delineation is even more pronounced. Furthermore, imprecise delineation may lead to sub-optimal treatment plans and unintended complications. Successful radiation therapy necessitates high geometric and dosimetric precision [41, 42, 43, 44, 29]. The rising trend of pencil beam scanning proton therapy has afforded us the ability to more accurately shape dose distributions to align with the treatment target, whilst minimizing the exposure to OARs [28, 45, 34, 46, 47, 48, 49, 50]. Nonetheless, the sharp dose gradients facilitated by these methods could potentially pose a risk if the delineation of structures is imprecise.Therefore, the rapid and precise delineation of OARs becomes even more vital for proton therapy, especially in regions where there is a dearth of experienced radiation oncologists.
@@
An automatic segmentation (AS) algorithm can potentially address these challenges, thereby considerably enhancing RT efficiency [51]. Numerous auto-segmentation methodologies have been proposed, including deformable image registration (DIR), atlas-based auto-segmentation, and the more recent deep learning-based segmentation (DLS) [52]. Although both DIR and atlas-based auto-segmentation have seen extensive implementation, their clinical utility is compromised due to limitations in accuracy and efficiency [53, 54]. The focus of auto-segmentation research has recently shifted towards AI methods, with a particular emphasis on those underpinned by deep learning (DL) [51, 55]. Over the past few years, the number of studies and clinical applications exploring DL-based segmentation of OARs in RT has proliferated, significantly improving the efficiency of auto-delineation in RT processes [56, 57, 58]. However, clinical feedback suggests a prominent issue with the current models: a lack of generalizability [59, 60]. These models often necessitate the training of a unique model for each RT site of one specific institution with medical images generated by a certain imaging protocol based on one specific imaging machine. Given the variety of RT sites, imaging protocols, and machines used, a multitude of independent auto-segmentation models need to be trained even within one institution, each requiring a considerable investment of time and a large volume of site/protocol/machine-specific training data. Some data harmonization methods have been proposed to mitigate this generalization problem with some limited success [61]. In practical clinical scenarios, the determination of the site to be delineated often requires the invocation of a site/protocol/machine-specific model. The complexity inherent in such scenarios, including mixed-site images and images from rare sites, can significantly degrade the precision of existing auto-segmentation models, and in some cases, cause errors. Thus, the development of an universal auto-delineation model, capable of handling multiple sites simultaneously, could significantly augment the precision and efficiency of auto-segmentation in RT processes, particularly in delineating OARs. An ideal solution would rely on limited data, generalize across protocols, modalities, anatomies, and institutions, and minimize human efforts. SAM demonstrates such potential but its performance on medical images remains unclear, especially for radiotherapy cases [62, 17].
@@
We evaluate SAM’s ability to perform zero-shot segmentation of medical images from multiple anatomical sites from clinical radiation oncology with CT images (prostate, lung, gastrointestinal, and head&neck). We assess SAM in "segment everything" mode where it generates masks for all objects and "box prompt" mode where users indicate regions of interest. Dice coefficient and Jaccard index are used to measure the spatial overlap between SAM’s predictions and ground truth clinical delineations.
@@
Our results provide a broad analysis of SAM for radiation oncology. SAM shows promise but performance varies significantly based on the disease sites and modalities. SAM excels at segmenting large, well-defined organs given unambiguous prompts but struggles with complex anatomies with ambiguous prompts, especially organs with unclear boundaries. Performance metrics suggest general trends but fuller qualitative analysis is warranted to determine practical utility, and there is much room for adapting and improving SAM for radiation oncology applications.
@@ 2.1 Segment Anything Model
The Segment Anything project proposed by Meta is a groundbreaking initiative that aims to democratize the world of image segmentation, a vital task in computer vision [18]. It consists of two key components: a substantial dataset for image segmentation, and the SAM, a promtable foundation model. Taking inspiration from the world of NLP, the project creates a vast dataset and a segmentation model, both open-sourced, thus opening up enormous possibilities. SAM, which can run in real-time in browsers, offers a highly automated image segmentation approach that requires minimal human intervention. It’s a deep learning model, trained on over 1 billion masks in 11 million images, which can cut out almost anything from an image. Unlike traditional models that require specialized training, SAM is generalizable and can respond to user prompts about specific areas to segment.
This model includes three essential components: an image encoder, a prompt encoder, and a mask decoder. Input images pass through an image encoder to produce an embedding, and the model can accept prompts as points, boxes, or rough masks. For more nuanced prompting, the authors are working on a version of SAM that accepts text input, similar to language models. The resultant segmented image comes with multiple valid masks and a confidence score, signifying the segmentation’s accuracy. SAM’s capabilities could revolutionize various fields, from augmenting reality through precise object identification to biomedical image segmentation, aiding cell microscopy analysis [63]. It can also be integrated with diffusion-based image generation models for efficient image inpainting and be utilized for generating semantic segmentation datasets, making it an exciting development in the realm of AI and computer vision [64].
@@ 2.2 SAM for Medical Imaging
Since SAM was proposed as an innovative framework for image segmentation, its application within the realm of medical imaging has increasingly become a topic of interest within the field [17] [65]. SAM has been tested under the ’everything’ mode for segmenting lesion regions across a range of anatomical structures, such as the brain, lung, and liver, and imaging modalities like CT and Magnetic Resonance Imaging (MRI). The experimental findings suggest that while SAM exhibits relative proficiency in segmenting organs with distinct boundaries, it encounters difficulties when trying to accurately identify lesions with irregular shapes and contours [66]. Another study drew comparisons between the SAM and the Brain Extraction Tool (BET) from the FMRIB Software Library in brain extraction tasks, and the quantitative analysis revealed superior segmentation performance by SAM compared to BET, thus illustrating SAM’s promising potential for use in brain segmentation tasks [67]. SAM was also applied to digital pathology segmentation tasks, which encompassed the segmentation of tumor and non-tumor tissue, as well as cell nuclei, on whole-slide images. The findings indicated that SAM exhibits excellent performance when segmenting large, interconnected objects. However, its reliability wavers when tasked with dense instance object segmentation, even when provided with all the target boxes [68]. SAM for polyp segmentation tasks using five benchmark datasets, all under the ’everything’ setting has been reported. The findings indicate that while SAM has the capacity to accurately segment polyps in certain instances, it significantly lags behind the performance of leading-edge methods, suggesting substantial room for improvement [69]. Extensive experimentation on multiple public datasets shows that the zero-shot segmentation capabilities of SAM are not sufficiently robust for direct application in medical image segmentation [17] [70] [63].
@@ 2.3 Cancer Segmentation
The field of cancer segmentation in medical imaging has seen considerable advancements in recent years, motivated by the potential to improve the accuracy of cancer detection, diagnosis, and radiotherapy. The goal of cancer segmentation is to accurately delineate the boundaries of cancerous regions within medical images.
@@
The recent explosion of deep learning techniques has revolutionized the field of cancer segmentation [74]. For instance, U-Net [75], a type of convolutional neural networks (a widely used architecture in medical image analysis [76, 77, 78, 79, 80]), has been applied to diverse tasks [81, 82, 39], including biomedical image segmentation such as lung cancer segmentation from CT scans and brain tumor segmentation from MRI images [75, 83, 81]. Developing fully automated systems for cancer segmentation is a trend in recent research [84]. These systems aim to minimize the need for manual intervention, thus reducing the time taken for diagnosis and the potential for human error.
The use of SAM for cancer segmentation has seen a significant increase since it was proposed. With point or bounding box prompts, SAM achieves competitive results comparable to supervised training models like U-net in the areas of liver tumor, breast tumor, and colon polyp segmentation [85, 86, 87]. By fine-tuning the image encoder, prompt encoder, and mask decoder using the same labeled data, SAM can even outperform state-of-the-art results in skin cancer segmentation [88]. These studies demonstrate that the pre-trained SAM is efficient and effective for computer-aided rapid cancer detection and diagnosis. Furthermore, it can also serve as a versatile foundation model for developing domain-specific models fine-tuned for targeted cancer segmentation.
@@ 3.1 Datasets
This study has received approval from the Institutional Review Board (IRB) at Mayo Clinic. Figure 1 illustrates the overall framework of our work to set up SAM for clinical case segmentation. Algorithm 1 further illustrates the internal algorithm process of SAM. We collected case images from the four most common sites in clinical radiotherapy, namely the prostate, gastrointestinal, lungs, and head & neck. We gathered 20 patients from each site, totaling 80 patients. From the perspective of clinical radiotherapy delineation, we divided the cases into two groups: the simple group includes the prostate and lungs, while the difficult group encompasses the gastrointestinal and head & neck. As SAM currently only supports 2D delineation, for a fair comparison, we evaluated the delineation of five typical 2D slices extracted from each patient that best represented the anatomical information. Based on the OARs that require attention during the formulation of the radiotherapy plan, we selected regions of interest for delineation comparison for different disease sites, following the recommendations of the Radiation Therapy Oncology Group (RTOG).
For the prostate, we selected the following regions of interest for evaluation: prostate, bladder, left femoral head, right femoral head, and rectum. For the lungs, we evaluated the following areas: left lung, right lung, heart, spinal cord, and esophagus. For the gastrointestinal, we assessed the liver, stomach, left kidney, right kidney, spinal cord, large bowel, and small bowel. For the head & neck, we evaluated the brain, left parotid, right parotid, spinal cord, mandible, left cochlea, and right cochlea. All manual delineations were performed by highly experienced radiation oncologists and meet the RTOG delineation standards, and are used in the formulation of clinical radiotherapy plans.